OpenStudy (anonymous):

Does the following sequence converge or diverge? If it converges then find the limit. An = ln(n) / ln(n+1).

OpenStudy (turingtest):

do you know l'Hospital's rule yet?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

taking the limit gives 0/0 so divergence test doesnt work

OpenStudy (turingtest):

what is the derivative of the numerator? what is the derivative of the denominator?

OpenStudy (anonymous):

derivative of ln(n) = 1/n and derivative of ln(n+1) = 1/(n+1)

OpenStudy (anonymous):

so if I take the limit of both they go to 0 right?

OpenStudy (anonymous):

wait this is a sequence so I don't use any of the tests

OpenStudy (anonymous):

so the sequence converges to 0?

OpenStudy (turingtest):

they don't to zero, one is in the denominator

OpenStudy (turingtest):

\[{\frac d{dn}\ln n\over\frac d{dn}\ln(n+1)}=?\]

OpenStudy (anonymous):

ohh right I didnt flip the denominator x.x so it's n+1/n which means it diverges to infinity.

OpenStudy (turingtest):

that does not diverge